write the equaton of a parabola that has a focus at (-1,7), has a minimum and the length from focus to vertex is 2 units
Answer by Edwin McCravy(20059) 
 write the equaton of a parabola that has a focus at (-1,7), has a minimum and the length from focus to vertex is 2 units
Since it has a minimum, it is of the form
(x-h)2 = 4a(y-k)
and the vertex is below the focus, and therefore opens upward.
Since the length from focus to vertex is 2 units,
1. the vertex is 2 units below the focus (-1,7), so
the vertex is (h,k) = (-1,5)
and
2. |a| = the distance from focus to vertex = 2, positive since
the parabola opens upward.
So the equation
(x-h)2 = 4a(y-k)
becomes
(x-(-1))2 = 4(2)(y-(5))
(x+1)2 = 8(y-5)
Edwin
|
|
|